OpenStudy (help_people):

The tenth, eleventh, and twelfth terms of a sequence are shown in the table below: Term number 10 11 12 Term -34 -38 -42 Which of the following shows the first five terms of the sequence? -2, 2, 6, 10, 14 2, -2, -6, -10, -14 -2, -6, -10, -14, -18 2, 6, 10, 14, 18

OpenStudy (welshfella):

note that to get from one term to the next you subtract 4

OpenStudy (help_people):

?

OpenStudy (welshfella):

-34 -4 = -38 -38-4 = -42

OpenStudy (welshfella):

this is an arithmetic sequence where the common difference is -4

OpenStudy (help_people):

ok

OpenStudy (welshfella):

so what do you think the 9th term is ? you subtract 4 to get the 10th so we can write -34 + 4 = a9 where a9 is the 9th term

OpenStudy (help_people):

-30

OpenStudy (welshfella):

right now you can keep on adding 4 until you get to the 5the term and see if it matches one of the options

OpenStudy (help_people):

a

OpenStudy (help_people):

is the answer

OpenStudy (welshfella):

no- you are guessing 8th term = -30 + 4 = -26 7th term = 26 + 4 = -22 - can you continue?

OpenStudy (help_people):

no I'm not i did kt

OpenStudy (welshfella):

ok

OpenStudy (help_people):

-22+4=-18 -18+4=-14 -14+4=-10 -10+4=-6 -6+4=-2 -2+4=2

OpenStudy (help_people):

see

OpenStudy (welshfella):

right sorry!

OpenStudy (help_people):

:D i chose the wrong one(my mistake) it looks like c but is missing 2

OpenStudy (welshfella):

you got there its B

OpenStudy (help_people):

thank you

OpenStudy (help_people):

may you help me w. another?

OpenStudy (help_people):

@welshfella

OpenStudy (welshfella):

ok